Table 11. Megans salary for three consecutive years, along with other values, are presented in the table below. Year 2011 012 2013 Salary Consumer Price Index Real Interest Rate $65,000$72,000$76,000 226 230 235 2.5 percent 2.7 percent 1.8 percent 12. Refer to Table 11. Suppose the consumer price index for 2013 is not necessarily 235. Then Megans 2013 salary represents more purchasing power than her 2011 salary as long as the consumer price index for 2013 is a. greater than 264.25. b. less than 264.25. c. greater than 258.71. d. less than 258.71. 13. Refer to Table 11. The nominal interest rate for 2012 is a.0.93 percent. b. 6.70 percent. c.4.44 percent. d. 4.47 percent.
Answer #1

Ans12) Megan salary 2010 shows more purchasing power than 2011 salary

$76000/cpi2013 >$65000/226

2013 cpi <$76000×226 /65000$

264. 25

Ans 13) nominal interest rate means refer interest rate before t inflation taking into account.

Nominal interest rates =real interest rates +inflation rate

Inflation rate =230-226/226*100=1.76

=2.7+1.76

=4.44%